The stockholders’ equity section of the balance sheet for Mann Equipment Co. at December 31, 2018, is as follows.

Stockholders Equity Paid-in capital Preferred stock,? par value, 4% cumulative, 290,000 shares authorized, 59,000 shares $590,006 1ssued and outstanding Common stock, $15 stated value, 340,000 shares authorized, 59,000?? shares issued and 885,000 49,000 outstanding Paid-in capital in excess of par-Preferred Paid-in capital in excess of stated 295,000 value-Common Total paid-in capital Retained earnings Treasury stock, 3,000 shares Total stockholders equity 1,819,000 440,000 42,000 $2,217,000 Note: The market value per share of the common stock is $38, and the market value per share of the preferred stock is $31
Answer #1

Share outstanding after the split = (59000-3000)*2 = 112000 Shares

Amount transferred from retained earnings = $0

Market price of common Stock after split = 38/2 = $19 per share